# Repeating nuclear transients from repeating partial tidal disruption events: reproducing ASASSN-14ko and AT2020vdq

### v1

@@ Abstract
Some electromagnetic outbursts from the nuclei of distant galaxies have been found to repeat on months-to-years timescales, and each of these sources can putatively arise from the accretion flares generated through the repeated tidal stripping of a star on a bound orbit about a supermassive black hole (SMBH), i.e., a repeating partial tidal disruption event (rpTDE). Here we test the rpTDE model through analytical estimates and hydrodynamical simulations of the interaction between a range of stars, which differ from one another in mass and age, and an SMBH. We show that higher-mass ($\gtrsim 1M_{\odot}$), evolved stars can survive many ($\gtrsim 10-100$) encounters with an SMBH while simultaneously losing $few\times 0.01M_{\odot}$, resulting in accretion flares that are approximately evenly spaced in time with nearly the same amplitude, quantitatively reproducing ASASSN-14ko. We also show that the energy imparted to the star via tides can lead to a change in its orbital period that is comparable to the observed decay in the recurrence time of ASASSN-14ko’s flares, $\dot{P}\simeq-0.0016$. Contrarily, lower-mass and less-evolved stars lose progressively more mass and produce brighter accretion flares on subsequent encounters for the same pericenter distances, leading to the rapid destruction of the star and cessation of flares. Such systems cannot reproduce ASASSN-14ko-like transients, but are promising candidates for recreating events such as AT2020vdq, which displayed a second and much brighter outburst compared to the first. Our results imply that the lightcurves of repeating transients are tightly coupled with stellar type.
@@ 1 Introduction
The standard TDE lightcurve is predicted to rise, peak, and decay monotonically with time $t$, and while the initial rise and peak are dependent on a number of factors involving the structure of the star (e.g., Lodato et al. 2009; Guillochon & Ramirez-Ruiz 2013; Golightly et al. 2019a; Law-Smith et al. 2020; Jankovič & Gomboc 2023), the asymptotic decline of the accretion rate with time scales either as $\propto t^{-5/3}$ if the disruption is complete (Rees, 1988; Phinney, 1989) or as $\propto t^{-9/4}$ if the disruption is partial (Coughlin & Nixon, 2019). Recently, however, there have been TDE flares [^1] Quasi-periodic eruptions (QPEs; e.g., Miniutti et al. 2019; Arcodia et al. 2021; Chakraborty et al. 2021; Quintin et al. 2023; Arcodia et al. 2024; Pasham et al. 2024) are qualitatively similar in terms of their highly cyclic behavior but are exclusively seen in the X-ray and have significantly shorter timescales (but see Evans et al. 2023; Guolo et al. 2024), and while there may be similarities in their physical origin that are related to TDEs, we restrict our phenomenological focus to longer-duration repeating events. – including ASASSN-14ko (Payne et al., 2021), AT2018fyk (Wevers et al., 2023), eRASSt-J045650 (Liu et al., 2024), and AT2020vdq (Somalwar et al., 2023) – that rebrighten on timescales of months to years following the initial peak, thus challenging this classic picture. One interpretation is that these sources represent repeating partial TDEs (rpTDEs), in which a star is orbiting an SMBH on a highly eccentric orbit, the pericenter distance of which is comparable to the partial disruption radius of the star; the latter is $\sim 2r_{\rm t}$, where $r_{\rm t}=R_{\star}\left(M_{\bullet}/M_{\star}\right)^{1/3}$ is the standard tidal disruption radius (Hills, 1975) with $R_{\star}$ the stellar radius and $M_{\bullet}$ and $M_{\star}$ the SMBH and stellar mass, respectively.
@@
Among the few rpTDE candidates, the longest-known and most heavily studied candidate is ASASSN-14ko, which has flared $\gtrsim 20$ times since its initial detection (Holoien et al., 2017; Kochanek et al., 2017; Payne et al., 2021). The energetics of the outbursts are consistent with the stripping of a star in which $\sim few\times 0.01M_{\odot}$ of mass is liberated – and subsequently accreted – during each pericenter passage if the luminosity is $L=0.1\dot{M}c^{2}$ (Cufari et al., 2022a). If ASASSN-14ko formed approximately contemporaneously with the first detection, meaning that there were no outbursts prior to the original detection (which must be the case if the star is near-solar, as it will only survive $\lesssim 100$ encounters if the per-encounter stripped mass is $\sim 0.01M_{\odot}$), then the orbital period of the star of $\sim 115$ days can be explained if the star was tidally captured from a tight binary (Cufari et al., 2022a). If the star was present for many more orbital periods prior to the original detection, then there are a number of mechanisms, including gravitational-wave inspiral and the interaction with a disc (Linial & Quataert, 2024), that can explain its current orbital period, though another mechanism must then be responsible for perturbing its pericenter distance to within $\sim 2r_{\rm t}$.
@@ 2 Toy Problem
However, the notion that the star is spun up repeatedly is generally not correct [^2] A secular increase in the rotational velocity could be achieved if the frequency of one of the stellar eigenmodes occurs in resonance with the orbital frequency, but it is difficult to see how changes in the stellar structure (induced by the rotation itself) would not cause the star to move off this resonance after only a few repeated encounters., because the star needs to be rotating at a rate that is smaller than the rotational velocity at pericenter in order to be efficiently torqued by the SMBH. Once the star achieves an angular velocity that is comparable to $\Omega_{\rm p}\simeq\sqrt{GM_{\bullet}}/r_{\rm p}^{3/2}$, where $r_{\rm p}\simeq few\times r_{\rm t}$ is the pericenter distance, the SMBH appears stationary (near pericenter) in the corotating frame of the star. In this case, then, there is no relative motion between the stellar surface and the speed with which the SMBH moves in the local frame of the star, thus preventing successive and efficient gravitational torques on the stellar body.
@@
Figure 2: Left: The ratio of the orbital period between pericenter passages $N$ and $N-1$ to the Keplerian orbital period, e.g., the value appropriate to $N=2$ represents the time taken to go from the first to second pericenter passage. Right: The ratio of the pericenter distance on the $N^{th}$ pericenter passage to the Keplerian pericenter distance. The values in these figures are consistent with the analytical expressions derived here; see Equation (2).
@@
For TDEs with $r_{\rm p}\simeq few\times r_{\rm t}=few\times R_{\star}\left(M_{\bullet}/M_{\star}\right)^{1/3}$, the relative changes in these quantities is of the order $\left(M_{\bullet}/M_{\star}\right)^{-2/3}\ll 1$. For the specific case considered here, with $e=0.9$ and $r_{\rm p}=300R_{\star}$, we have $\Delta T/T\simeq 6.6\times 10^{-4}$ and $\Delta r_{\rm p}/r_{\rm p}\simeq 2.2\times 10^{-5}$. These numbers are in good agreement with Figure 2, the left panel of which shows the ratio of the orbital period between pericenter passages $N$ and $N-1$, $T_{\rm N,N-1}$ to the Keplerian orbital period, $T$, while the right panel shows the ratio of the pericenter distance, $r_{\rm p,N}$, to the Keplerian pericenter distance $r_{\rm p}$.
@@
This toy problem will overestimate the angular momentum (and angular velocity) imparted to the star in the limit that $r_{\rm p}\gg r_{\rm t}$, the reason being that the quadrupole moment of the “dumbbell” is effectively the maximum possible value that a star could have. Nonetheless, it is illustrative in demonstrating that a star can only be given an angular velocity comparable to $\sim\Omega_{\rm p}$, implying that – provided $r_{\rm p}$ is outside the tidal radius – a star is capable of surviving for many repeated encounters with an SMBH, thus powering the repeated flares observed in ASASSN-14ko.
Additionally, Equation (2) can estimate the period decay rate of rpTDEs. For ASASSN-14ko specifically, a period of $T=114$ days and a SMBH mass of $10^{7}M_{\odot}$ – as inferred by Payne et al. (2023) – yields a semimajor axis of $a\simeq 1.5\times 10^{15}$ cm. If the pericenter distance is comparable to the tidal radius, then $r_{\rm p}\simeq 2\times r_{\rm t}=2R_{\star}\left(M_{\bullet}/M_{\star}\right)^{1/3}$, and the eccentricity of the orbit satisfies [^3] It is interesting to note that the mass of the black hole drops out of this expression.
@@
$\frac{\Delta T}{T}\simeq\frac{3}{4}\left(1-e\right)^{-1}\left(\frac{M_{\bullet}}{M_{\star}}\right)^{-2/3}\simeq 0.001.$ (5)
@@ 3.1 Simulation Setup
To simulate the disruption process, the relaxed star was placed at an initial distance of $5r_{\rm t}$ from the SMBH, with the COM on a parabolic Keplerian orbit. Upon passing through pericenter, the tidal field of the SMBH partially disrupts the star and strips off a fraction of its outer envelope. Approximately half of the stripped debris is gravitationally bound to the SMBH and accretes onto it. The surviving core in a partial TDE reduces the hydrodynamical timestep of the simulations relative to the dynamical time of the disrupted debris. Thus, to calculate the fallback rate of bound stellar debris, the surviving stellar core was replaced with a point mass, roughly 2 days after the COM reached pericenter, at which point the core had receded to a distance of $>15r_{\rm t}$ from the SMBH (for low mass stars, having smaller tidal radii, this distance is larger) as was done in, e.g., Miles et al. (2020); Golightly et al. (2019b). The fallback rate is calculated as the rate at which particles from the disrupted debris stream return to the accretion radius of the SMBH, which is defined as the inner $3r_{\rm t}$ of the computational domain for our simulations. The fallback rate of particles from the disrupted debris stream closely tracks the accretion rate onto the SMBH, subject to the assumption of efficient disc circularization and negligible viscous delays (see Mockler et al. 2019; Nicholl et al. 2022 for observational evidence that suggests that viscous delays are small).
@@
Simulating the subsequent encounters by evolving the star on its highly eccentric orbit is computationally intractable, since the orbital time of these events, which ranges from months-to-years, amounts to thousands of dynamical times of the disrupted star. To surmount this issue, we simulated each encounter by allowing the COM of the star to evolve on its parabolic orbit around the SMBH for $\sim 2$ days past its pericenter, which amounts to $\gtrsim 100$ dynamical times of the star (but still orders of magnitude shorter than the orbital time). The surviving core was then translated back to the initial position of the original star (equal to $5r_{\rm t}$) by calculating and then subtracting the COM velocity and position from every SPH particle, and subsequently adding back (to each particle in the core) those of the original stellar orbit. Figure 3 shows the density profile of the partially disrupted star for a $3M_{\odot}$ TAMS star, at a time when the COM has evolved for $\sim 2$ days past its pericenter. As seen in the figure, there is a distinct dichotomy between the core and stream particles, with the exception of a few particles that are situated in the intermediate region. For each simulation, we chose a density cutoff that lies between the tidally disrupted debris stream and the dense central structure that we identify as the core. The core, comprised of particles having a density greater than the threshold value, was translated back on its orbit to simulate the subsequent encounters. We verified that once a distinct core had formed, the density profile remained effectively unmodified, and thus, performing the core-translation routine a day earlier (or later) does not have a measurable effect on the fallback rate. The fallback rate for each successive orbit was obtained by replacing the core with a sink particle using the procedure described above, and tracking the rate of return of particles from the debris stream to the SMBH.

### v2

@@ Abstract
Some electromagnetic outbursts from the nuclei of distant galaxies have been found to repeat on months-to-years timescales, and each of these sources can putatively arise from the accretion flares generated through the repeated tidal stripping of a star on a bound orbit about a supermassive black hole (SMBH), i.e., a repeating partial tidal disruption event (rpTDE). Here we test the rpTDE model through analytical estimates and hydrodynamical simulations of the interaction between a range of stars, which differ from one another in mass and age, and an SMBH. We show that higher-mass ($\gtrsim 1M_{\odot}$), evolved stars can survive many ($\gtrsim 10-100$) encounters with an SMBH while simultaneously losing $few\times 0.01M_{\odot}$, resulting in accretion flares that are approximately evenly spaced in time with nearly the same amplitude, quantitatively reproducing ASASSN-14ko. We also show that the energy imparted to the star via tides can lead to a change in its orbital period that is comparable to the observed decay in the recurrence time of ASASSN-14ko’s flares, $\dot{P}\simeq-0.0026$. Contrarily, lower-mass and less-evolved stars lose progressively more mass and produce brighter accretion flares on subsequent encounters for the same pericenter distances, leading to the rapid destruction of the star and cessation of flares. Such systems cannot reproduce ASASSN-14ko-like transients, but are promising candidates for recreating events such as AT2020vdq, which displayed a second and much brighter outburst compared to the first. Our results imply that the lightcurves of repeating transients are tightly coupled with stellar type.
@@ 1 Introduction
The standard TDE lightcurve is predicted to rise, peak, and decay monotonically with time $t$, and while the initial rise and peak are dependent on a number of factors involving the structure of the star (e.g., Lodato et al. 2009; Guillochon & Ramirez-Ruiz 2013; Golightly et al. 2019a; Law-Smith et al. 2020; Jankovič & Gomboc 2023), the asymptotic decline of the accretion rate with time scales either as $\propto t^{-5/3}$ if the disruption is complete (Rees, 1988; Phinney, 1989) or as $\propto t^{-9/4}$ if the disruption is partial (Coughlin & Nixon, 2019). Recently, however, there have been TDE flares [^1] Quasi-periodic eruptions (QPEs; e.g., Miniutti et al. 2019; Giustini et al. 2020; Arcodia et al. 2021; Chakraborty et al. 2021; Quintin et al. 2023; Arcodia et al. 2024; Pasham et al. 2024b) are qualitatively similar in terms of their highly cyclic behavior but are exclusively seen in the X-ray and have significantly shorter timescales (but see Evans et al. 2023; Guolo et al. 2024), and while there may be similarities in their physical origin that are related to TDEs, we restrict our phenomenological focus to longer-duration repeating events. – including ASASSN-14ko (Payne et al., 2021), AT2018fyk (Wevers et al., 2023), eRASSt-J045650 (Liu et al., 2024b), and AT2020vdq (Somalwar et al., 2023) – that rebrighten on timescales of months to years following the initial peak, thus challenging this classic picture. One interpretation is that these sources represent repeating partial TDEs (rpTDEs), in which a star is orbiting an SMBH on a highly eccentric orbit, the pericenter distance of which is comparable to the partial disruption radius of the star; the latter is $\sim 2r_{\rm t}$, where $r_{\rm t}=R_{\star}\left(M_{\bullet}/M_{\star}\right)^{1/3}$ is the standard tidal disruption radius (Hills, 1975) with $R_{\star}$ the stellar radius and $M_{\bullet}$ and $M_{\star}$ the SMBH and stellar mass, respectively.
@@
Among the few rpTDE candidates, the longest-known and most heavily studied candidate is ASASSN-14ko, which has flared $\gtrsim 20$ times since its initial detection (Holoien et al., 2017; Kochanek et al., 2017; Payne et al., 2021). The energetics of the outbursts are consistent with the stripping of a star in which $\sim few\times 0.01M_{\odot}$ of mass is liberated – and subsequently accreted – during each pericenter passage if the luminosity is $L=0.1\dot{M}c^{2}$ (Cufari et al., 2022a). If ASASSN-14ko formed approximately contemporaneously with the first detection, meaning that there were no outbursts prior to the original detection (which must be the case if the star is near-solar, as it will only survive $\lesssim 100$ encounters if the per-encounter stripped mass is $\sim 0.01M_{\odot}$), then the orbital period of the star of $\sim 115$ days can be explained if the star was tidally captured from a tight binary (Cufari et al., 2022a). However, if the star loses $\sim 0.001M_{\odot}$ on each encounter, with the mass stripped per-encounter increasing gradually with time, then the system could have existed for many more orbital periods prior to the original detection, and there are a number of mechanisms, including gravitational-wave inspiral and the interaction with a disc (Linial & Quataert, 2024), that can then explain its current orbital period, though another mechanism must then be responsible for perturbing its pericenter distance to within $\sim 2r_{\rm t}$.
@@ 2 Toy Problem
However, the notion that the star is spun up repeatedly is generally not correct [^2] A secular increase in the rotational velocity could be achieved if the frequency of one of the stellar eigenmodes occurs in resonance with the orbital frequency, but it is difficult to see how changes in the stellar structure (induced by the rotation itself) would not cause the star to move off this resonance after only a few repeated encounters., because the star needs to be rotating at a rate that is smaller than the rotational velocity at pericenter in order to be efficiently torqued by the SMBH. Once the star achieves an angular velocity that is comparable to $\Omega_{\rm p}\simeq\sqrt{GM_{\bullet}\left(1+e\right)}/r_{\rm p}^{3/2}$, where $r_{\rm p}\simeq few\times r_{\rm t}$ is the pericenter distance, the SMBH appears stationary (near pericenter) in the corotating frame of the star. In this case, then, there is no relative motion between the stellar surface and the speed with which the SMBH moves in the local frame of the star, thus preventing successive and efficient gravitational torques on the stellar body.
@@
Figure 2: Left: The relative difference of the time between pericenter passages $N$ and $N-1$ to the Keplerian orbital period, e.g., the value appropriate to $N=2$ represents the time between the first and second pericenter passage, minus and subsequently divided by the Keplerian orbital period $T$. Right: The relative difference between the pericenter distance on the $N^{\rm th}$ pericenter passage to the Keplerian pericenter distance $r_{\rm p}$. The values in these figures are consistent with the analytical expressions derived here; see Equation (2). These data are appropriate to the orange curve in Figure 1, from which it is apparent that the largest changes in the rotational velocity of the dumbbell correspond to the largest changes in both the orbital period and the pericenter, which is expected.
@@
For TDEs on bound orbits with $r_{\rm p}\simeq few\times r_{\rm t}=few\times R_{\star}\left(M_{\bullet}/M_{\star}\right)^{1/3}$, the relative changes in these quantities scale with the mass ratio as $\left(M_{\bullet}/M_{\star}\right)^{-2/3}\ll 1$. The contribution from the factor $(1+e)/(1-e)$ becomes increasingly significant in the limit $e\rightarrow 1$. For the specific case considered here, with $e=0.9$ and $r_{\rm p}=300R_{\star}$, we have $\Delta T/T\simeq 3.2\times 10^{-4}$ and $\Delta r_{\rm p}/r_{\rm p}\simeq 2.2\times 10^{-5}$. These numbers are in good agreement with Figure 2, the left panel of which shows the relative difference of the time between pericenter passages $N$ and $N-1$ to the Keplerian orbital period $T$, $\Delta T_{\rm N,N-1}$, e.g., the value at $N=2$ gives the time between the second and first pericenter passage, minus and divided by the Keplerian orbital period. The right panel shows the relative difference between the pericenter distance of the Nth pericenter passage and the Keplerian pericenter distance. These data are appropriate to the orange curve in Figure 1; note that – in agreement with expectations – the largest changes in the orbital period and pericenter correspond to the largest angular velocities of the dumbbell.
@@
This toy problem will overestimate the angular momentum imparted to the star in the limit that $r_{\rm p}\gg r_{\rm t}$, the reason being that the quadrupole moment of the “dumbbell” is permanent and maximal (compared to that of a star, which is itself excited by tides). Nonetheless, it is illustrative in demonstrating that a star can only be given an angular velocity comparable to $\sim\Omega_{\rm p}$, implying that – provided $r_{\rm p}$ is outside the tidal radius – a star is capable of surviving for many repeated encounters with an SMBH, thus powering the repeated flares observed in ASASSN-14ko.
Additionally, Equation (2) can estimate the period decay rate of rpTDEs. For ASASSN-14ko specifically, a period of $T=114$ days and an SMBH mass of $10^{7}M_{\odot}$ – as inferred by Payne et al. (2023) – yields a semimajor axis of $a\simeq 1.5\times 10^{15}$ cm. If the pericenter distance is comparable to the tidal radius, then $r_{\rm p}\simeq 2\times r_{\rm t}=2R_{\star}\left(M_{\bullet}/M_{\star}\right)^{1/3}$ (we note that high-mass and more evolved stars can give rise to partial TDEs and survive multiple encounters even when the pericenter distance lies closer to the tidal radius; Coughlin & Nixon 2022a), and the eccentricity of the orbit satisfies [^3] It is interesting to note that the mass of the black hole drops out of this expression.
@@
$\frac{\Delta T}{T}\simeq\frac{3}{2}\left(\frac{1+e}{1-e}\right)\left(\frac{M_{\bullet}}{M_{\star}}\right)^{-2/3}\simeq 0.003.$ (5)
@@ 3.1 Simulation Setup
To simulate the disruption process, the relaxed star was placed at an initial distance of $5r_{\rm t}$ from the SMBH, with the COM on a parabolic Keplerian orbit. Upon passing through pericenter, the tidal field of the SMBH partially disrupts the star and strips off a fraction of its outer envelope. Approximately half of the stripped debris is gravitationally bound to the SMBH and accretes onto it. The surviving core in a partial TDE reduces the hydrodynamical timestep of the simulations relative to the dynamical time of the disrupted debris. Thus, to calculate the fallback rate of bound stellar debris, the surviving stellar core was replaced with a point mass roughly 2 days after the COM reached pericenter, at which point the core had receded to a distance of $>15r_{\rm t}$ from the SMBH (for low mass stars, having smaller tidal radii, this distance is larger) as was done in, e.g., Miles et al. (2020); Golightly et al. (2019b). The fallback rate is calculated as the rate at which particles from the disrupted debris stream return to the accretion radius of the SMBH, which is defined as the inner $3r_{\rm t}$ of the computational domain for our simulations. The fallback rate of particles from the disrupted debris stream closely tracks the accretion rate onto the SMBH, subject to the assumption of efficient disc circularization and negligible viscous delays (see Mockler et al. 2019; Nicholl et al. 2022 for observational evidence that suggests that viscous delays are small).
Simulating the subsequent encounters by evolving the star on its highly eccentric orbit is computationally intractable, since the orbital time of these events, which ranges from months-to-years, amounts to thousands of dynamical times of the disrupted star. To surmount this issue, we simulated each encounter by allowing the COM of the star to evolve on its parabolic orbit around the SMBH for $\sim 2$ days past its pericenter, which amounts to $\gtrsim 100$ dynamical times of the star (but still orders of magnitude shorter than the orbital time). The surviving core was then translated back to the initial position of the original star (equal to $5r_{\rm t}$) by calculating and then subtracting the COM velocity and position from every SPH particle, and subsequently adding back (to each particle in the core) those of the original stellar orbit. Figure 3 shows the density profile of the partially disrupted star for a $3M_{\odot}$ TAMS star, at a time when the COM has evolved for $\sim 2$ days past its pericenter. As seen in the figure, there is a distinct dichotomy between the core and stream particles, with the exception of a few particles that are situated in the intermediate region. For each simulation, we chose a density cutoff that lies between the tidally disrupted debris stream and the dense central structure that we identify as the core. The core, comprised of particles having a density greater than the threshold value, was translated back on its orbit to simulate the subsequent encounters. We verified that once a distinct core had formed, the density profile remained effectively unmodified, implying that performing the core-translation routine a day earlier (or later) does not have a measurable effect on the fallback rate. The fallback rate for each successive orbit was obtained by replacing the core with a sink particle using the procedure described above, and tracking the rate of return of particles from the debris stream to the SMBH.
@@
3.2 Results for a $3M_{\odot}$ TAMS star
One of our primary goals in this work is to test the viability of the rpTDE model in providing an explanation for the repeated flares observed from ASASSN-14ko. To do this, we describe here the results of simulating the repeated partial disruption of a $3M_{\odot}$ TAMS star by a $10^{6}M_{\odot}$ SMBH, with an impact parameter $\beta=1$. The impact paramater $\beta\equiv r_{\rm t}/r_{\rm p}$ quantifies the strength of the encounter between the star and the SMBH. For a high-mass star at its late evolutionary stages, the analytical prediction for the critical impact parameter $\beta_{\rm c}$ required for its complete disruption exceeds this value by an appreciable amount (Coughlin & Nixon 2022a; from Figure 4 of Bandopadhyay et al. (2024) a $3M_{\odot}$ TAMS star has a critical impact parameter $\beta_{\rm c}\sim 6$, though we note that general relativistic effects may reduce this value somewhat for very massive SMBHs; e.g., Beloborodov et al. 1992; Kesden 2012; Gafton et al. 2015; Tejeda et al. 2017; Gafton & Rosswog 2019; Ryu et al. 2020; Jankovič & Gomboc 2023). Thus, $\beta=1$ results in only a small amount of mass lost, and the star should – subject to the influence of the imparted rotation discussed above – be able to survive for many encounters while simultaneously fueling accretion events.
Each successive pericenter passage spins up the tidally stripped core by a small fraction of its break-up velocity. Figure 4 shows the velocity of the fluid comprising the core in its COM coordinates (i.e., both COM position and velocity removed), which qualitatively demonstrates that the core is indeed rotating following its tidal interaction with the SMBH. In the top panel of Figure 5, we show the angular velocity $\Omega$ imparted to the core during the first fifteen pericenter passages as a function of cylindrical radius $s$ within the core. The primary $y-$axis shows the angular velocity normalized by the break-up angular velocity of the core, $\Omega_{\rm c}=\sqrt{GM_{\rm c}/R_{\rm c}^{3}}$. Here, $R_{\rm c}\approx 0.2R_{\odot}$ is the core radius for a $3M_{\odot}$ TAMS star, defined as the radial distance from the center of the star at which the self-gravitational field of the star is maximized (Coughlin & Nixon, 2022a), and $M_{\rm c}\approx 0.3M_{\odot}$ is the mass contained within the core. The secondary $y-$axis shows the angular velocity normalized by the angular velocity of the COM at pericenter, $\Omega_{\rm p}=\sqrt{2GM_{\bullet}/r_{\rm p}^{3}}$. This figure demonstrates that while the rotation rate is comparable to $\Omega_{\rm p}$, the imparted angular velocity is a small fraction of the break-up angular velocity of the core. The core is thus relatively unmodified, and this allows the bulk of the star to survive many encounters as it tidally interacts with the SMBH. The bottom panel of Figure 5 shows the corresponding linear velocity $v_{\varphi}\equiv s\Omega$, normalized by $\sqrt{GM_{\rm c}/R_{\rm c}}$.
Additionally, we note that in our simulations, the thermal energy that is dissipated viscously (numerically) is not put back into the gas, and is instead assumed to be efficiently lost from the system; this choice is motivated by the fact that the outer layers of the star can be subject to excessive numerical heating (see, e.g., Norman et al. 2021; Coughlin & Nixon 2022b). There should, however, be some level of physical heating in response to the dissipation of kinetic energy imparted by tides, and depending on where the thermal energy is deposited, it may instead be trapped within the star, leading to its more rapid destruction. To assess the impact of viscous heating and to demonstrate that it does not have a strong impact on our results and corresponding conclusions, in Appendix A we present a detailed investigation of the effects of both numerical resolution and the thermodynamic prescription.